OpenStudy (anonymous):

. A student has to secure 40% marks to pass. He gets 40 and fails by 40 marks. Find the maximum marks.

OpenStudy (nirmalnema):

passing marks is 80 because student got 40 and failed by 40 so basically you have to find 80 is the 40% of which no..? so 80/x * 100 = 40 8000/x=40 x=8000/40 x=200=maximum marks

OpenStudy (nirmalnema):

do you get it @Nixon

OpenStudy (anonymous):

Yes I got it thanx

OpenStudy (nirmalnema):

ok your welcome

OpenStudy (anonymous):

Can u answer 1 more question??

OpenStudy (nirmalnema):

yup sure i will try..

OpenStudy (anonymous):

7/3+35/18+54/20 = then X =?

OpenStudy (nirmalnema):

take the LCM and add them simply

OpenStudy (anonymous):

Yeah i know.. But when i search it up on yahoo there answer and method is totally different..

OpenStudy (nirmalnema):

what is your ans...?

OpenStudy (anonymous):

6.977

OpenStudy (nirmalnema):

yup its the correct ans its 1256/180 420+350+486 / 180

OpenStudy (anonymous):

Hmmm Thanks..